### v1

@@ Abstract
In this paper, we propose a novel class of Piecewise Deterministic Markov Processes (PDMP) that are designed to sample from constrained probability distributions $\pi$ supported on a convex set $\mathcal{M}$. This class of PDMPs adapts the concept of a mirror map from convex optimisation to address sampling problems. Such samplers provides unbiased algorithms that respect the constraints and, moreover, allow for exact subsampling. We demonstrate the advantages of these algorithms on a range of constrained sampling problems where the proposed algorithm outperforms state of the art stochastic differential equation-based methods.
@@ 1 Introduction
where $\mathcal{M}$ is a convex subset of $\mathbb{R}^{d}$ is central to Bayesian computation and arises in a number of applications such as image processing [1]), statistical mechanics ([2]), and inverse problems ([3]). Markov chain Monte Carlo (MCMC) methods deal with this problem by constructing a Markov chain that is stationary with respect to $\pi$. Two widely used approaches to construct such a Markov chain are as follows: 1) Based on an accept-reject mechanism [4]; 2) Based on stochastic differential equations (SDEs) such as the Langevin SDE [5]. Algorithms based on the first approach may suffer from poor mixing properties, especially in high dimensions [6]. On the other hand, SDE-based algorithms can exhibit faster convergence properties at the cost of some (controlled) bias [7, 8, 9].
@@
Piecewise Deterministic Markov Processes (PDMPs) were introduced in the early 1980s by [10] in the context of biological and communication systems. They are non-diffusive, irreversible Markov processes that evolve deterministically between random times. Similarly to the case of SDE samplers, it is possible to design PDMPs that are invariant with respect to $\pi$. This has led to variety of sampling algorithms such as the Zig-Zag sampler (ZZS), [11], the Bouncy Particle Sampler (BPS), [12], the Boomerang Sampler (BS), [13] and the Coordinate Sampler(CS), [14] to name a few. Implementation of these algorithms is discussed in [15].
@@
PDMPs are particularly attractive for MCMC due to three properties: i) they can be simulated exactly; ii) they are irreversible, thus potentially converging faster towards $\pi$ [16, 17], iii) they allow for unbiased estimation using stochastic gradients. This is useful in the context of large data sets where the cost of the algorithm can be reduced to $\mathcal{O}(1)$ cost per iteration as shown in [18].
@@
All the algorithms mentioned above deal with unconstrained sampling problems, that is $\mathcal{M}=\mathbb{R}^{d}$. When $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$, there are several approaches for SDE-based samplers that borrow ideas from convex optimisation. In particular, a natural extension of the projected gradient descent [19], the projected Langevin Monte Carlo method ([20]) incorporates a projection back to $\mathcal{M}$ after every Euler-Maruyama step. Alternatively, one can use a regularised version of the indicator function $\bm{1}_{\mathcal{M}}$ to define a probability density with support over $\mathbb{R}^{d}$ and employ different numerical discretizations of the underlying SDE leading for example to the Moreau Yosida Unadjusted Langevin Algorithm (MYULA) ([21, 22]) or the Forward Backward Unadjusted Langevin algorithm (FBULA) ([23]). The mirror Langevin algorithms use the mirror map to ensure that the samples produced reside in $\mathcal{M}$ [24, 25, 26]. Finally, [27, 28] add a Riemannian metric to the space $\mathcal{M}$ and use suitable discretizations of geometrically inspired SDEs.
@@
Ideally, one would like to combine the ability of the SDE-based samplers mentioned above to deal with the fact that $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$ with the desirable features of PDMPs, namely unbiasedness, and the ability to subsample. To the best of our knowledge, PDMP samplers have not been considered in this setting, and the central purpose of this paper is to provide a construction and implementation of such samplers. In particular, we introduce a framework that adapts PDMPs to include mirror maps and thus allows sampling in the case when $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$. In addition, we specify a more detailed choice of PDMP called mirror Zig-Zag sampler (MZZS) and establish conditions for its ergodicity.
@@
The remainder of this paper is organized as follows: in Section 2, we discuss in detail SDE-based samplers both in the constrained and the unconstrained case. Furthermore, in Section 3 we give a broad introduction to PDMPs in the unconstrained setting. Section 4 extends these ideas to the constrained setting and explains the role of the mirror map for PDMPs, while Section 5 shows a number of numerical experiments that illustrate the benefits of mirror PDMPs.
2 SDE based samplers
@@ 2.1 The unconstrained case
There exist several different SDEs that have $\pi$ as their invariant distribution. As shown in [29] these SDEs can be summarised by the following family
@@
for some semi-definite diffusion matrix $S(x)$ and skew-symmetric matrix $J(x)$. Here $W_{t}$ is a standard $d$-dimensional Brownian motion.
@@
The simplest method for solving (3) is the Euler-Maruyama method [30]
$X_{n+1}=X_{n}-\Delta t\nabla U(X_{n})+\sqrt{2\Delta t}\xi_{n},$ (4)
@@ 2.2 The constrained case
There are several different approaches for dealing with the case when $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$. In particular, the projected Langevin Monte Carlo method (PLMC) ([20])
$X_{n+1}=P_{\mathcal{M}}(X_{n}-\Delta t\nabla U(X_{n})+\sqrt{2\Delta t}\xi_{n})$ (5)
adds a projection $P_{\mathcal{M}}$ back to $\mathcal{M}$ after each Euler-Maruyama step (4) to ensure that iterates of (5) remain within $\mathcal{M}$. Alternatively, one can use a regularised version of the indicator function $\bm{1}_{\mathcal{M}}$ to construct the potential
@@
which now defines a probability density with support over the whole of $\mathbb{R}^{d}$. Discretizing, the Langevin equation (3) with the Euler-Maruyama method gives
$X_{n+1}=X_{n}-\Delta t\nabla U(X_{n})+\frac{\Delta t}{\varepsilon}\left(P_{\mathcal{M}}(X_{n})-X_{n}\right)+\sqrt{2\Delta t}\xi_{n}.$ (7)
One drawback of (7) relates to the fact that one needs to operate with a small value of the regularisation parameter $\varepsilon$. This results in $\Delta t$ being at most $\mathcal{O}(\varepsilon)$, which in turn leads to slow mixing. On the other hand (5) doesn’t impose an additional constraint on $\Delta t$ and can be shown to be ergodic for any value of $\Delta t$ provided $\mathcal{M}$ is bounded and connected. However, using large values of $\Delta t$ leads to a large bias.
@@ 2.2.1 SDEs exploiting the mirror map
An alternative way to deal with constraints similar to those already discussed is by using a mirror map. In the case of convex optimization [19] mirror descent
$\zeta_{n+1}$ $=\zeta_{n}-\Delta t\nabla U(x_{n}),$ (8a)
$x_{n+1}$ $=\nabla\psi^{*}(\zeta_{n+1}),$ (8b)
@@
A barrier function associated to the set $\mathcal{M}$ is a map $\psi:\mathcal{M}\to\mathbb{R}$ which is closed, proper, strictly convex, twice continuously differentiable and such that
$\|\nabla\psi({x_{i}})\|\to\infty\quad\text{as}\ i\to\infty,$
whenever $x_{1},x_{2},\ldots$ is a sequence in $\mathcal{M}$ converging to $x\in\partial\mathcal{M}$ where $\partial\mathcal{M}$ denotes the boundary of $\mathcal{M}$.
For such a barrier function $\psi$, the mirror map $\nabla\psi:\mathcal{M}\to\mathbb{R}^{d}$ is a bijection from the convex set $\mathcal{M}$ to $\mathbb{R}^{d}$ [31, Corollary 26.3.1]. Moreover, the inverse of $\nabla\psi$ is $\nabla\psi^{*}$ where
@@
refers to the Legendre-Fenchel conjugate of $\psi$ [31, Corollary 23.5.1]. We will refer to $\mathcal{M}$ as the primal space and the image of the mirror map as the dual space, under the assumptions of Definition 2.1 the dual space is $\mathbb{R}^{d}$. Note that (8) can be seen as a discretisation of the mirror gradient flow [32]
@@
As discussed in the unconstrained case, one can write many different SDEs that have the same invariant distribution. Thus, when utilising the idea of the mirror map for SDE-based sampling, [24] and [25, 26] proposed two different SDEs for constrained sampling. More precisely, [24] proposed
@@
where $V$ is the negative log density of the push-forward distribution $\tilde{\pi}=(\nabla\psi)_{\#}\pi$ and is given by
@@
The Euler discretization of (10) is given by
$\zeta_{n+1}$ $=\zeta_{n}-\Delta t\nabla V(\zeta_{n})+\sqrt{2\Delta t}\xi_{n},$ (12a)
$x_{n+1}$ $=\nabla\psi^{*}(\zeta_{n+1}),$ (12b)
and hereafter we will refer to this as the mirror Langevin algorithm with additive noise (MLAa).
@@
is proposed in [25, 26]. An Euler discretization of (13) was proposed in [25] and analysed in [33], while the following more sophisticated discretization was proposed in [26]
$x_{n+1/2}=\nabla\psi^{*}(\nabla\psi(x_{n})-\Delta t\nabla U(x_{n}))$ (14a)
$x_{n+1}=\nabla\psi^{*}(Z_{\Delta t}),\qquad\text{where}\quad\begin{cases}\mathrm{d}Z_{t}=\sqrt{2}[\nabla^{2}\psi^{*}(Z_{t})]^{-1/2}\mathrm{d}W_{t},\\ Z_{0}=\nabla\psi(x_{n+1/2}).\end{cases}$ (14b)
We will refer to this discretization as the mirror Langevin algorithm with multiplicative noise (MLAm). As in [26], we discretize the $Z_{t}$ dynamics in (14b) using 10 steps of Euler-Maruyama algorithm for each step of MLAm.
@@
$S(x)=[\nabla^{2}\psi(X_{t})]^{-2}=[\nabla^{2}\psi^{*}(\nabla\psi(X_{t}))]^{2}$
@@
$S(x)=[\nabla^{2}\psi(X_{t})]^{-1}=\nabla^{2}\psi^{*}(\nabla\psi(X_{t})).$

### v2

@@ Abstract
In this paper, we propose a novel class of Piecewise Deterministic Markov Processes (PDMPs) that are designed to sample from probability distributions $\pi$ supported on a convex set $\mathcal{M}$. This class of PDMPs adapts the concept of a mirror map from convex optimisation to address sampling problems. The corresponding algorithms provide unbiased samples that respect the constraints and, moreover, allow for exact subsampling. We demonstrate the advantages of these algorithms against a range of constrained sampling problems where the proposed algorithms outperform state of the art stochastic differential equation-based methods.
@@ 1 Introduction
is central to Bayesian computation, and arises in a number of applications such as image processing [green2015bayesian], statistical mechanics [Leimkuhler15], and inverse problems [Stuart_2010]. Here $U:\mathcal{M}\to\mathbb{R}$ is a given function, and $\mathcal{M}$ is a convex subset of $\mathbb{R}^{d}$. Markov chain Monte Carlo (MCMC) methods deal with this problem by constructing a Markov chain that is stationary with respect to $\pi$. Two widely used approaches to construct such a Markov chain are as follows: 1) Based on an accept-reject mechanism [metropolis1953equation]; 2) Based on stochastic differential equations (SDEs) such as the Langevin SDE [Roberts96]. Algorithms based on the first approach may suffer from poor mixing properties, especially in high dimensions [DM19]. On the other hand, SDE-based algorithms can exhibit faster convergence properties at the cost of some (controlled) bias [durmus2017nonasymptotic, JMLR:v22:21-0453, Dalalyan19].
@@
Piecewise Deterministic Markov Processes (PDMPs) were introduced in the early 1980s by [davis1984piecewise] in the context of biological and communication systems. They are non-diffusive, irreversible Markov processes that evolve deterministically between random times. Similarly to the case of SDE samplers, it is possible to design PDMPs that are invariant with respect to $\pi$. This has led to variety of sampling algorithms such as the Zig-Zag sampler (ZZS), [bierkens2019ergodicity], the Bouncy Particle Sampler (BPS), [bouchard2018bouncy], the Boomerang Sampler (BS), [bierkens2020boomerang] and the Coordinate Sampler(CS), [wu2020coordinate] to name a few. Implementation of these algorithms is discussed in [fearnhead2018piecewise].
@@
PDMPs are particularly attractive for MCMC due to three properties: i) they can be simulated exactly under suitable conditions. [^1] A detailed discussion about challenges with exact simulation is given in Section 3.2.; ii) they are irreversible, thus potentially converging faster towards $\pi$ [diaconis2000analysis, bierkens2016non], iii) they allow for unbiased estimation using stochastic gradients. This is useful in the context of large data sets where the cost of the algorithm can be reduced to $\mathcal{O}(1)$ cost per iteration as shown in [bierkens2019zig].
@@
All the algorithms mentioned above deal with unconstrained sampling problems, that is $\mathcal{M}=\mathbb{R}^{d}$. When $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$, there are several approaches for SDE-based samplers that borrow ideas from convex optimisation. In particular, a natural extension of the projected gradient descent [Beck17], the projected Langevin Monte Carlo method [NIPS2015_c0f168ce] incorporates a projection back to $\mathcal{M}$ after every Euler-Maruyama step. Alternatively, one can use a regularised version of the indicator function $\bm{1}_{\mathcal{M}}$ to define a probability density with support over $\mathbb{R}^{d}$ and employ different numerical discretizations of the underlying SDE leading for example to the Moreau Yosida Unadjusted Langevin Algorithm (MYULA) [durmus2018efficient, brosse2017sampling] or the Forward Backward Unadjusted Langevin algorithm (FBULA) [eftekhari2023forward]. The mirror Langevin algorithms use the mirror map to ensure that the samples produced reside in $\mathcal{M}$ [hsieh2018mirrored, zhang20a, Chewi20]. Finally, [GC11, BLS25] add a Riemannian metric to the space $\mathcal{M}$ and use suitable discretizations of geometrically inspired SDEs.
@@
Ideally, one would like to combine the ability of the SDE-based samplers mentioned above to deal with the fact that $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$ with the desirable features of PDMPs, namely unbiasedness, and the ability to subsample. One approach [bierkens2023methods] to do this is to view the probability distribution $\pi$ as a discontinuous function of $\mathbb{R}^{d}$ and then add reflections to standard PDMPs to construct a PDMP that respects the constraints. In this paper we take a different approach and introduce a framework that adapts PDMPs to include mirror maps and thus allows sampling in the case when $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$. In addition, we specify a more detailed choice of PDMP called mirror Zig-Zag sampler (MZZS) and establish conditions for its ergodicity.
@@
The remainder of this paper is organized as follows: in Section 2, we discuss in detail SDE-based samplers both in the constrained and the unconstrained cases. Furthermore, in Section 3 we give a broad introduction to PDMPs in the unconstrained setting. Section 4 extends these ideas to the constrained setting and explains the role of the mirror map for PDMPs, while Section 5 shows a number of numerical experiments that illustrate the benefits of mirror PDMPs.
2 SDE-based samplers
@@ 2.1 The unconstrained case
There exist several different SDEs that have $\pi$ as their invariant distribution. As shown in [Fox15] these SDEs can be summarised by the following family
@@
for some semi-definite diffusion matrix $S(x)$ and skew-symmetric matrix $J(x)$. Here $(W_{t})_{t\geq 0}$ is a standard $d$-dimensional Brownian motion.
@@
The simplest method for solving (3) is the Euler-Maruyama method [KP92]
$\bar{X}_{n+1}=\bar{X}_{n}-\Delta t\nabla U(\bar{X}_{n})+\sqrt{2\Delta t}\xi_{n},$ (4)
@@ 2.2 The constrained case
There are several different approaches for dealing with the case when $\mathcal{M}$ is a proper convex subset of $\mathbb{R}^{d}$. One such approach is based on the reflected SDE (RSDE)
$dX_{t}=-\nabla U(X_{t})dt+\sqrt{2}dW_{t}-n(X_{t})dL_{t},$
where $L_{t}$ is the boundary local time (a non decreasing process that increases only when $X_{t}\in\partial\mathcal{M}$) and $n$ is the outward unit normal vector field. There is a variety of discretisation methods applicable to this RSDE see for example [cattiaux2017invariant, Bossy04, leimkuhler2020simplest, pettersson1997penalization, Gobet2001]. Here, we will concentrate on two discretisations: the projected Langevin Monte Carlo method (PLMC) [NIPS2015_c0f168ce] and the penalty method which corresponds to MYULA for this RSDE. The projected Langevin Monte Carlo method (PLMC) given by
$\bar{X}_{n+1}=P_{\mathcal{M}}(\bar{X}_{n}-\Delta t\nabla U(\bar{X}_{n})+\sqrt{2\Delta t}\xi_{n})$ (5)
adds a projection $P_{\mathcal{M}}$ back to $\mathcal{M}$ after each Euler-Maruyama step (4) to ensure that iterates of (5) remain within $\mathcal{M}$. The penalty method [pettersson1997penalization] uses a regularised version of the indicator function $\bm{1}_{\mathcal{M}}$ to construct the potential
@@
which now defines a probability density with support over the whole of $\mathbb{R}^{d}$ where $\|\cdot\|$ denotes the Euclidean norm. Discretizing, the Langevin equation (3) with the Euler-Maruyama method for the potential $U^{\varepsilon}$ gives
$\bar{X}_{n+1}=\bar{X}_{n}-\Delta t\nabla U(\bar{X}_{n})+\frac{\Delta t}{\varepsilon}\left(P_{\mathcal{M}}(\bar{X}_{n})-\bar{X}_{n}\right)+\sqrt{2\Delta t}\xi_{n}.$ (7)
One drawback of (7) relates to the fact that one needs to operate with a small value of the regularisation parameter $\varepsilon$. This results in $\Delta t$ being at most $\mathcal{O}(\varepsilon)$, which in turn leads to slow mixing. On the other hand (5) does not impose an additional constraint on $\Delta t$ and can be shown to be ergodic for any value of $\Delta t$ provided $\mathcal{M}$ is bounded and connected. However, using large values of $\Delta t$ leads to a large bias.
Note there are other approaches such as those based on reflective Hamiltonian Monte Carlo [chalkis2021truncated] which consider the case where $\mathcal{M}$ is a convex polytope and Wall Hamiltonian Monte Carlo [pakman2014exact] which is specialised to the case where $U$ is a quadratic function.
@@ 2.2.1 SDEs exploiting the mirror map
An alternative way to deal with constraints similar to those already discussed is by using a mirror map. In the case of convex optimization [Beck17] mirror descent
$\bar{\zeta}_{n+1}$ $=\bar{\zeta}_{n}-\Delta t\nabla U(\bar{X}_{n}),$ (8a)
$\bar{X}_{n+1}$ $=\nabla\psi^{*}(\bar{\zeta}_{n+1}),$ (8b)
@@
A barrier function associated to the set $\mathcal{M}$ is a map $\psi:\text{int}(\mathcal{M})\to\mathbb{R}$ which is, proper, closed, strictly convex, twice continuously differentiable and such that
$\|\nabla\psi({x_{n}})\|\to\infty\quad\text{as}\ n\to\infty,$
whenever $x_{1},x_{2},\ldots$ is a sequence in $\mathcal{M}$ converging to $x\in\partial\mathcal{M}$ where $\partial\mathcal{M}$ denotes the boundary of $\mathcal{M}$ and $\text{int}(\mathcal{M})$ its interior.
Some examples of commonly used barrier functions [hsieh2018mirrored, zhang20a, srinivasan2024fast] along with the corresponding set $\mathcal{M}$ are presented in Table 1.
Domain $\mathcal{M}$ Barrier function
Rectangle: $\mathcal{M}=\{x\in\mathbb{R}^{d}:a_{i}\leq x_{i}\leq b_{i}\}$ $\psi(x)=-\sum_{i=1}^{d}\log(x_{i}-a_{i})+\log(b_{i}-x_{i})$,
Euclidean ball: $\mathcal{M}=\{x\in\mathbb{R}^{d}:\|x\|\leq 1\}$ $\psi(x)=-\log(1-\|x\|)-\|x\|$
$\text{Polytope}(A,b):\mathcal{M}=\{x\in\mathbb{R}^{d}:Ax\leq b\},\ A\in\mathbb{R}^{m\times d},\ b\in\mathbb{R}^{m}$ $\psi(x)=-\sum_{i=1}^{m}\log(b_{i}-a_{i}^{\top}x)$
$\text{Lorentz cone:}\ \mathcal{M}=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<x_{2}\}$ $\psi(x)=-\log(x_{2}^{2}-x_{1}^{2})$
$\text{Positive orthant: }\ \mathcal{M}=(a_{i},+\infty)^{d}=\{x\in\mathbb{R}^{d}:x_{i}>a_{i}\}$ $\psi(x)=\frac{1}{2}\|x\|^{2}-\sum_{i=1}^{d}\log(x_{i}-a_{i})$
$\text{Simplex}:\Delta_{d}=\{x\in\mathbb{R}^{d}:\sum_{i=1}^{d}x_{i}=1\},\ x_{d}=1-\sum_{i=1}^{d-1}x_{i}$ $\psi(x)=\sum_{i=1}^{d-1}x_{i}\log(x_{i})+x_{d}\log(x_{d})$
Table 1: Examples of convex sets and their corresponding barrier functions.
For a given barrier function $\psi$, the mirror map $\nabla\psi:\mathcal{M}\to\mathbb{R}^{d}$ is a bijection from the convex set $\mathcal{M}$ to $\mathbb{R}^{d}$ [tyrrell1970convex, Corollary 26.3.1]. Moreover, the inverse of $\nabla\psi$ is $\nabla\psi^{*}$ where
@@
refers to the Legendre-Fenchel conjugate of $\psi$ [tyrrell1970convex, Corollary 23.5.1] in which $x^{\top}$ denotes the transpose of the vector $x$. We will refer to $\mathcal{M}$ as the primal space and the image of the mirror map as the dual space. Under the assumptions of Definition 2.1 the dual space is $\mathbb{R}^{d}$ [tyrrell1970convex, Theorem 26.5]. Note that (8) can be seen as a discretisation of the mirror gradient flow [NY83]
@@
As discussed in the unconstrained case, one can write many different SDEs that have the same invariant distribution. Thus, when utilising the idea of the mirror map for SDE-based sampling, [hsieh2018mirrored, zhang20a, Chewi20] proposed two different SDEs for constrained sampling. More precisely, [hsieh2018mirrored] proposed
@@
where $V$ is the negative log of the unnormalized density of the push-forward distribution $\tilde{\pi}=(\nabla\psi)_{\#}\pi$ and is given by
@@
Here $\circ$ denotes function composition and $\det$ is the determinant. The Euler-Maruyama discretization of (10) is given by
$\bar{\zeta}_{n+1}$ $=\bar{\zeta}_{n}-\Delta t\nabla V(\bar{\zeta}_{n})+\sqrt{2\Delta t}\xi_{n},$ (12a)
$\bar{X}_{n+1}$ $=\nabla\psi^{*}(\bar{\zeta}_{n+1}),$ (12b)
and hereafter we will refer to this as the mirror Langevin algorithm with additive noise (MLAa). A Metropolis adjusted version of this algorithm was considered and analysed in [zhang20a, srinivasan2024fast], which we will refer to as the Metropolis Adjusted Mirror Langevin Algorithm (MAMLA).
@@
is proposed in [zhang20a, Chewi20]. An Euler discretization of (13) was proposed in [zhang20a] and analysed in [li2022mirror], while the following more sophisticated discretization was proposed in [Chewi20]
$\bar{X}_{n+1/2}=\nabla\psi^{*}(\nabla\psi(\bar{X}_{n})-\Delta t\nabla U(\bar{X}_{n}))$ (14a)
$\bar{X}_{n+1}=\nabla\psi^{*}(Z_{\Delta t}),\qquad\text{where}\quad\begin{cases}dZ_{t}=\sqrt{2}[\nabla^{2}\psi^{*}(Z_{t})]^{-1/2}dW_{t},\\ Z_{0}=\nabla\psi(\bar{X}_{n+1/2}).\end{cases}$ (14b)
We will refer to this discretization as the mirror Langevin algorithm with multiplicative noise (MLAm). As in [Chewi20], we discretize the $Z_{t}$ dynamics in (14b) using 10 steps of Euler-Maruyama algorithm for each step of MLAm.
@@
$S(x)=[\nabla^{2}\psi({x})]^{-2}=[\nabla^{2}\psi^{*}(\nabla\psi({x}))]^{2}$
@@
$S(x)=[\nabla^{2}\psi({x})]^{-1}=\nabla^{2}\psi^{*}(\nabla\psi({x})).$ (15)
Remark 2.1.
One can view the mirror map $\psi$ as imposing a Riemannian manifold structure on the set $\mathcal{M}$. The Riemannian metric on this manifold is given by $\nabla^{2}\psi$. With this structure one can construct the Riemannian manifold Langevin SDE as in [GC11], this is the SDE (2.1) with $S$ given by (15) and hence (13) coincides with the Riemannian manifold Langevin SDE with the metric $\nabla^{2}\psi$, see [zhang20a] for further details.